# Test-Time Scaling of Reasoning Models for Machine Translation

### v1

@@ Abstract
Zihao Li,1 Shaoxiong Ji,2,3 Jörg Tiedemann1 1University of Helsinki  2 University of Turku  3 ELLIS Institute Finland firstname.lastname@{1helsinki.fi, 2utu.fi}
@@ 1 Introduction
Large language models (LLMs) have dramatically advanced machine translation (MT), evolving from statistical and neural paradigms to systems capable of handling diverse languages, domains, and complexities with unprecedented accuracy (Lyu et al., 2023; Kocmi et al., 2024; Zhu et al., 2024; Cui et al., 2025). Recent developments in Reasoning Models (RMs)—models designed to incorporate structured reasoning processes like Chain-of-Thought (CoT)—have further transformed MT by reframing it as a cognitive task requiring contextual analysis, cultural adaptation, and self-reflection (Liu et al., 2025). For instance, RMs can resolve ambiguities in stylized texts and maintain coherence across documents, thereby outperforming traditional LLMs in semantically demanding scenarios (Ye et al., 2025).
@@
This paper investigates these open questions through three research questions:
@@
  • RQ3: Does test-time scaling work in post-editing? We evaluate TTS in self-correction scenarios, assessing its role in refining initial translations.
@@
  • We show that the effectiveness of TTS is unlocked by domain-specific fine-tuning, which aligns the model’s reasoning process with task requirements. For DRT models fine-tuned on specific domain data, performance improves with budget on in-domain tasks and saturates once models naturally stop increasing their internal token usage, suggesting an emergent alignment between optimal reasoning depth and task demands. This alignment largely disappears out of domain.
@@ 2.2 Models
  • Cogito: Two models, sized 3B and 8B, were included (Deep Cogito, 2025). Cogito-3B and Cogito-8B are trained on top of Llama-3.2-3B and Llama-3.1-8B (Dubey et al., 2024), respectively, and similarly implement hybrid reasoning capabilities with controllable switching between generation and reasoning modes.
  • DRT: Three models from this family were evaluated. These models are fine-tuned from existing LLMs using the training set of MetaphorTrans (Wang et al., 2025a). Specifically, DRT-7B, DRT-8B, and DRT-14B are built upon Qwen2.5-7B-Instruct (Qwen et al., 2024), Llama-3.1-8B-Instruct, and Qwen2.5-14B-Instruct respectively.
@@ 2.5 Post-Editing
Figure 2: Average GRB scores of Qwen-3 and Cogito models across all datasets with varying thinking budgets.
@@ 3.1 Effectiveness of Test-Time Scaling
Our initial investigation focused on the efficacy of test-time scaling for general-purpose RMs, including the Qwen-3 and Cogito families, as well as the proprietary Grok-3-Mini model. These models were evaluated “out-of-the-box" without any fine-tuning on our benchmark datasets. Figure 2 plots the average GRB scores across all datasets against varying thinking-token budgets for the Qwen-3 and Cogito model series. After a small initial performance gain when moving from a zero budget to a minimal budget (e.g., 100 tokens), the models’ performance curves almost completely plateau in most cases. Further increases in the budget up to 2000 tokens yield no meaningful improvements, indicating that simply allocating more computational steps does not enable the models to produce more refined or accurate translations when they lack specific task-related knowledge.
@@
On the in-domain MetaphorTrans task, we observe a clear and consistent positive correlation between the thinking budget and translation performance. As the thinking budget increases from 100 to around 500 tokens, both the number of generated tokens and the GEA scores steadily rise.
@@
This efficient alignment vanishes on the other out-of-domain literary translation tasks. The most striking counterexample is the document-level WMT24-Literary task. Here, the actual thinking tokens continue to scale almost linearly with the budget, indicating the models are using the provided extra capacity to “think" longer. Yet, this extended reasoning does not translate into better performance; the GEA scores remain erratic and show no consistent improvement. This disconnect suggests the models are engaged in unproductive or unfocused reasoning, “spinning their wheels" without the specialized knowledge required for this different type of literary translation. This dichotomy underscores our central argument: test-time scaling is most effective when fine-tuning has equipped a model with not only domain-specific knowledge but also an efficient strategy for how and when to apply it.
@@ 4 Related Work
LLMs have rapidly advanced machine translation, with growing emphasis on open-weight systems that balance translation quality, multilingual coverage, and general-purpose capabilities. Seed-X (Cheng et al., 2025) introduces a 7B-parameter open-source family of translation-oriented LLMs trained on large-scale monolingual and bilingual corpora across 28 languages, achieving performance competitive with closed-source systems such as GPT-4o (Hurst et al., 2024) and Gemini-2.5 (Comanici et al., 2025). Similarly, Hunyuan-MT (Zheng et al., 2025) develops 7B-parameter models: Hunyuan-MT and Hunyuan-MT-Chimera, the latter integrates multiple outputs under a “slow thinking” paradigm to yield higher-quality translations, and ranks first in the WMT2025 shared task across 30 of 31 directions. Tower+ (Rei et al., 2025) addresses the trade-off between translation specialization and general-purpose ability by combining continued pretraining, supervised fine-tuning, preference optimization, and reinforcement learning with verifiable rewards.

### v2

@@ Abstract
Zihao Li,1 Shaoxiong Ji,2,3,1 Jörg Tiedemann1 1University of Helsinki  2 ELLIS Institute Finland  3 University of Turku firstname.lastname@{1helsinki.fi, 3utu.fi}
@@ 1 Introduction
Large language models (LLMs) have dramatically advanced machine translation (MT), evolving from statistical and neural paradigms to systems capable of handling diverse languages, domains, and complexities with unprecedented accuracy (Lyu et al., 2023; Kocmi et al., 2024; Zhu et al., 2024; Cui et al., 2025; Hendy et al., 2023). Recent developments in Reasoning Models (RMs), models designed to incorporate structured reasoning processes like Chain-of-Thought (CoT), have further transformed MT by reframing it as a cognitive task requiring contextual analysis, cultural adaptation, and self-reflection (Liu et al., 2025). For instance, RMs can resolve ambiguities in stylized texts and maintain coherence across documents, thereby outperforming traditional LLMs in semantically demanding scenarios (Ye et al., 2025).
@@
This paper investigates these open questions by distinguishing between two TTS workflows: Direct Translation (single-pass CoT scaling) and Post-Editing (compute-scaled self-correction). We structure our investigation through three research questions:
@@
  • RQ3: Does test-time scaling work in post-editing? We evaluate TTS in self-correction scenarios, assessing its role in refining initial translations when being allocated with specific compute budgets.
@@
  • We show that the effectiveness of TTS is unlocked by domain-specific fine-tuning, which aligns the model’s reasoning process with task requirements. For DRT models fine-tuned on specific domain data, performance improves with budget on in-domain tasks and saturates once models naturally stop increasing their internal token usage, suggesting an emergent alignment between optimal reasoning depth and task demands. This alignment largely disappears out of the domain.
@@ 2.2 Models
  • Cogito: Two models, sized 3B and 8B, were included (Deep Cogito, 2025). Cogito-3B and Cogito-8B are trained on top of Llama-3.2-3B and Llama-3.1-8B (Grattafiori et al., 2024), respectively, and similarly implement hybrid reasoning capabilities with controllable switching between generation and reasoning modes.
  • DRT: Three models from this family were evaluated. These models are fine-tuned from existing LLMs using the training set of MetaphorTrans (Wang et al., 2025a). Specifically, DRT-7B, DRT-8B, and DRT-14B are built upon Qwen2.5-7B-Instruct (Yang et al., 2024), Llama-3.1-8B-Instruct, and Qwen2.5-14B-Instruct respectively.
@@ 2.5 Post-Editing
Figure 2: Average GRB scores and average actual thinking tokens of Qwen-3 and Cogito models across all datasets with varying thinking budgets.
@@ 3.1 Effectiveness of Test-Time Scaling
Our initial investigation focused on the efficacy of test-time scaling for general-purpose RMs, including the Qwen-3 and Cogito families, as well as the proprietary Grok-3-Mini model. These models were evaluated “out-of-the-box" without any fine-tuning on our benchmark datasets. Figure 2 plots the average GRB scores (right axis) alongside the actual thinking tokens generated (left axis) for the Qwen-3 and Cogito model series. After a small initial performance gain when moving from a zero budget to a minimal budget (e.g., 100 tokens), the models’ performance curves almost completely plateau in most cases. A critical observation emerges from the thinking-token curves: despite being allocated budgets up to 2000 tokens, most general-purpose models fail to utilize this capacity. The actual reasoning length typically saturates around 600 tokens. This suggests that these models reach a “reasoning ceiling” where they autonomously terminate the deliberation process, indicating that simply allocating more computational steps does not enable the models to produce more refined or accurate translations when they lack specific task-related knowledge.
@@
On the in-domain MetaphorTrans task, we observe a clear and consistent positive correlation between the thinking budget and translation performance. We consider the model’s natural stopping point, where the ‘Actual tokens’ curve plateaus around 500 tokens, as the realistic baseline. As the thinking budget increases from 100 to this limit, both the number of generated tokens and the GEA scores steadily rise. This monotonic improvement indicates that the model is performing valid, necessary reasoning steps.
@@
This efficient alignment vanishes on the other out-of-domain literary translation tasks. The most striking counterexample is the document-level WMT24-Literary task. Here, the actual thinking tokens continue to scale almost linearly with the budget, indicating the models are using the provided extra capacity to “think" longer. [^1] We attribute this distinct behavior to text granularity: unlike the out-of-domain paragraph-level tasks (LitEval-Corpus) where limited context leads models to exhaust reasoning paths and saturate early, the extensive context in document-level translation allows models to continuously expand their reasoning loops to consume the available budget. Yet, this extended reasoning does not translate into better performance; the GEA scores remain erratic and show no consistent improvement. This disconnect suggests the models are engaged in unproductive or unfocused reasoning, “spinning their wheels" without the specialized knowledge required for this different type of literary translation. This dichotomy underscores our central argument: test-time scaling is most effective when fine-tuning has equipped a model with not only domain-specific knowledge but also an efficient strategy for how and when to apply it.
@@ 4 Related Work
The application of LLMs to machine translation has witnessed a paradigm shift. Foundational research demonstrated that general-purpose models, such as GPT-3, possess remarkable few-shot translation capabilities, challenging traditional supervised systems (Brown et al., 2020). Subsequent empirical studies systematically evaluated these capabilities, revealing that while LLMs excel in high-resource languages and specific domains, they often require careful adaptation to match state-of-the-art baselines (Hendy et al., 2023; Zhu et al., 2024). To address these limitations, researchers have developed sophisticated in-context learning strategies (Agrawal et al., 2023) and diverse prompting techniques (Vilar et al., 2023). Notably, recent work by Wu et al. (2025) highlights the efficacy of simple re-translation prompts over complex reasoning in general LLMs, suggesting that iterative refinement can significantly boost performance without elaborate chains of thought. Concurrently, the field is moving towards open-weight models that balance broad multilingual competence with translation specialization. Seed-X (Cheng et al., 2025) introduces a 7B-parameter open-source family of translation-oriented LLMs trained on large-scale monolingual and bilingual corpora across 28 languages, achieving performance competitive with closed-source systems such as GPT-4o (Hurst et al., 2024) and Gemini-2.5 (Comanici et al., 2025). Similarly, Hunyuan-MT (Zheng et al., 2025) develops 7B-parameter models: Hunyuan-MT and Hunyuan-MT-Chimera, the latter integrates multiple outputs under a “slow thinking” paradigm to yield higher-quality translations, and ranks first in the WMT2025 shared task across 30 of 31 directions (Kocmi et al., 2025). Tower+ (Rei et al., 2025) addresses the trade-off between translation specialization and general-purpose ability by combining continued pretraining, supervised fine-tuning, preference optimization, and reinforcement learning with verifiable rewards.
@@ 5 Conclusion
Acknowledgments
The work has received funding from the Digital Europe Programme under grant agreement No 101195233 (OpenEuroLLM). The authors wish to acknowledge CSC - IT Center for Science, Finland, and LUMI supercomputers, owned by the EuroHPC Joint Undertaking, for providing computational resources. Shaoxiong Ji received support from the CA21167 COST action UniDive, funded by COST (European Cooperation in Science and Technology).